# Augmenting Subspace Optimization Methods with Linear Bandits

### v1

@@ 1 Introduction
$\min_{x\in\mathbb{R}^{d}}f(x).$ (1)
We make the following assumption on $f$, which will hold throughout this manuscript.
@@
  • $f:\mathbb{R}^{d}\to\mathbb{R}$ is continuously differentiable on a level set $L(x_{0}):=\{z\in\mathbb{R}^{d}:f(z)\leq f(x_{0})\}$.
  • $f$ is bounded below on $L(x_{0})$; that is, there exists $x_{*}\in\mathbb{R}^{d}$ such that $f(x_{*})\leq f(x)$ for all $x\in L(x_{0})$.
  • $\nabla f$ is globally Lipschitz with constant $L$ on the level set $L(x_{0})$; that is, $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|$ for all $x,y\in L(x_{0})$.
@@
In some areas of computational science, the computation of derivatives of objective function $f$ may be prohibitive even when 1 is satisfied. This is traditionally the scope of derivative-free optimization (DFO) methods; see, for example, [3, 16, 31]. Alternatively, for such functions $f$ that are provided as computer codes, algorithmic differentiation (AD) can provide numerically accurate derivative computations. A celebrated result (see, e.g., [22][Section 3.2]) proves that, in the absence of constraints on memory, reverse-mode AD can compute a numerical gradient of a function in a small factor times the wall-clock time of performing a zeroth-order function evaluation. Of course, these memory assumptions are unrealistic; and in practice, the computation of derivatives via reverse-mode AD is challenging, if not prohibitive. For a more thorough discussion of such challenges, see, for example, [26][Section 1]. In situations where reverse-mode AD is too difficult, however, forward-mode AD is sometimes tractable. Informally, forward-mode AD computes directional derivatives (in potentially arbitrary, that is, not coordinate-aligned, directions). The cost of a single directional derivative in this mode is approximately the cost of a single zeroth-order evaluation of $f$. This is, of course, an informal statement; see [26] for practical insight into more precise quantification of the true costs of forward-mode AD. Thus, one may view forward-mode AD as an oracle for computing directional derivatives (in arbitrary directions) at a cost similar to that of performing finite difference approximations of derivatives but with generally superior accuracy.
In computational settings that preclude analytic derivatives and efficient reverse-mode AD, subspace optimization methods are of particular interest for the solution of (1) and have seen recent theoretical growth. For the sake of exposition, we broadly define subspace methods as any iterative method of minimization that can be described by the informal statement provided in Framework 1.
1 Input: Initial point $x_{0}\in\mathbb{R}^{d}$, number of iterations $K$, function $f$ satisfying 1.
@@
3       Choose a subspace dimension $p_{k}\leq d$, and choose a matrix $S_{k}\in\mathbb{R}^{d\times p_{k}}$.
4       Construct some model, $m_{k}:\mathbb{R}^{p_{k}}\to\mathbb{R}$, such that $m_{k}(z)$ approximates $f(x_{k}+S_{k}z)$ for all $z\in\mathbb{R}^{p_{k}}$ in a neighborhood of $0\in\mathbb{R}^{p_{k}}$.
5       Approximately minimize $m_{k}$ in a neighborhood of $0\in\mathbb{R}^{p_{k}}$ to obtain $z_{*}\in\mathbb{R}^{p_{k}}$.
6       Update the incumbent $x_{k+1}\leftarrow x_{k}+S_{k}z_{*}$, provided $f(x_{k}+S_{k}z_{*})$ yields sufficient decrease.
@@
Framework 1 Informal definition of a subspace optimization method
@@
Deterministic subspace methods have been studied for decades; Krylov methods, and more specifically nonlinear conjugate gradient methods, fit within Framework 1 by implicitly choosing $S_{k}$ to define a (typically) two-dimensional subspace based on past gradient evaluations. (L)BFGS methods implicitly generate $S_{k}$ such that the span of the columns of $S_{k}$ is equal to the span of a (limited) memory of gradients and past incumbent/gradient displacements. Block coordinate descent methods can be viewed as choosing $S_{k}$ as a (potentially deterministic) subset of columns of a $d$-dimensional identity matrix. The former two types of subspace methods require the computation of (full-space) gradients, so these methods are not within the scope of our computational setting. Block coordinate descent methods do fit in our computational setting but are not the topic of this paper.
@@
Recent developments have seen the application of randomized matrix sketching [32, 50, 34, 37] as a means to choose $S_{k}$ within algorithms in Framework 1. In a computational setting that permits forward-mode AD, the local model $m_{k}$ of Framework 1 could be simply a first-order Taylor model in the affine subspace defined by $x_{k}$ and $S_{k}$. Computing $S_{k}^{\top}\nabla f(x_{k})$ amounts to computing a (relatively small) number of directional derivatives in the randomized directions given in the columns of $S_{k}$. In the derivative-free setting, an arbitrary model $m_{k}$ is constructed in the (relatively small) subspace defined by $x_{k}$ and $S_{k}$; $m_{k}$ might exhibit the property that the zeroth- and first-order errors made by the model are comparable to those made by a first-order Taylor model. In either the forward-mode AD setting or the derivative-free setting, and as we will expound upon, the theoretical success of randomized subspace methods hinges upon some notion that the randomly selected affine subspace in each iteration exhibits, with sufficiently high probability, a sufficiently large normalized inner product with the gradient at the incumbent; that is, roughly speaking,
$\|(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top}\nabla f(x_{k})\|\approx\|\nabla f(x_{k})\|.$ (2)
@@
Providing a guarantee like (2) typically relies on some variant of the Johnson–Lindenstrauss lemma [27], a surprising result that demonstrates, roughly speaking, that in high dimensions, a low-dimensional subspace defined by $S_{k}$ (with $p_{k}$ independent of $d$) drawn from an appropriate distribution achieves a property resembling (2) with high probability.
@@ 1.1 Contributions
In this manuscript we propose a novel deterministic method for choosing subspaces $S_{k}$ within subspace optimization methods like those described by Framework 1. Instead of directly employing randomized sketches in an effort to (probabilistically) align the random subspace with a gradient, we demonstrate how this alignment problem can be described through the lens of linear bandit methods in a nonstationary environment. Because 1 guarantees that the gradient of $f$ is Lipschitz continuous, we treat the incumbent update in a generic subspace optimization method as an exogenous process resulting in bounded drifts in a gradient.
@@
More specifically, we adapt a linear upper confidence bound (UCB) bandit method [4, 2] to choose each unit-length one-dimensional sketch $S_{k}$ so that $S_{k}$ is maximally aligned with $\nabla f(x_{k})$; that is, we aim to maximize $|(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top}\nabla f(x_{k})|$ in every iteration. The linear UCB method maintains a confidence ellipsoid describing a belief on the dynamically changing value of $\nabla f(x_{k})$ by fitting a linear regression model using a history of observed (approximations to) $S_{k}^{\top}\nabla f(x_{k})$. The maximizer over this confidence ellipsoid of an inner product with a unit-norm vector is chosen on each iteration as the one-dimensional sketch $S_{k}$. Such a UCB method ensures a balance between exploration of subspaces—the focus of randomized sketching approaches to subspace optimization methods in the literature—and exploitation of subspaces that have recently been observed to exhibit large gradients. The linear UCB method is introduced and analyzed in Section 3.
@@
In Section 3 we also demonstrate sublinear dynamic regret [7, 8] of this linear UCB method in terms of its ability to approximate the dynamically changing gradient $\nabla f(x_{k})$. Because of our choice to view the optimization dynamics as an exogenous process affecting the gradient $\nabla f(x_{k})$, we stress that this is not a regret bound for the optimization algorithm. However, our regret analysis will elucidate how the selection of a particular subspace optimization algorithm will affect the regret incurred by the linear UCB method. As an illustration, we will consider perhaps the simplest possible optimization algorithm that could fit into Framework 1—a subspace variant of fixed stepsize gradient descent that employs randomized sketching—to make these connections clear. In particular, given a sequence of matrices $\{S_{k}\}$, define the associated sequence of embedding matrices $\{P_{k}:=S_{k}(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top}\}\subset\mathbb{R}^{d\times d}$, and let $\{\alpha_{k}\}$ denote a strictly positive sequence of stepsizes. With these choices made, we define a subspace gradient descent method via the iteration
$x_{k+1}\leftarrow T(x_{k};S_{k})\triangleq x_{k}-\alpha_{k}P_{k}\nabla f(x_{k}).$ (3)
@@
We are particularly concerned in this manuscript with practicalities; and so, motivated by various pieces of the analysis in Section 3, we propose practical extensions in Section 4. We conclude this paper by proposing and testing our linear UCB mechanism in an algorithm that can be viewed as a specific instance of a quadratic regularization method with randomized sketching proposed in [9] in a setting where we simulate the computation of forward-mode AD gradients; the method is essentially that described by (3) but with a dynamically adjusted stepsize $\alpha_{k}$. In the derivative-free setting, we propose a simple extension of the DFO method POUNDers [49] that employs our linear UCB mechanism to select subspaces.
@@ 1.2 Related Work
The theoretical and practical development of methods that have the general form of Framework 1 has enjoyed much recent attention. Matrix sketching [32, 50, 34, 37], which entails the random generation of $S_{k}$ in Framework 1 from some distribution on matrices, has emerged as a theoretical tool for analyzing such methods, often supported by the powerful Johnson–Lindenstrauss lemma [27]. In [29] the authors propose a method in Framework 1 that looks much like the iteration (3). The authors prove convergence of the method under standard assumptions about convexity of $f$; in a follow-up paper [30], the same authors considered the same iteration but with $\nabla f(x_{k})$ approximated by finite differences. The works [21, 52, 19] and [24, 58] propose and analyze randomized Newton methods and stochastic subspace cubic Newton methods, respectively, that extend the iteration (3) by also using second-derivative information projected onto subspaces defined by $S_{k}$. Pilanci and Wainwright [41] also similarly consider randomized Newton methods in the case where the objective Hessian admits a matrix square root everywhere on its domain. Cartis et al. [9] consider imposing various globalization strategies, such as trust-region and quadratic regularization, into algorithms that fall into Framework 1; the authors also specialize their methods to nonlinear least-squares objectives. Miyashi et al. [35] consider a quasi-Newton approximation of projected Hessian information to be used with projected gradients $S_{k}^{\top}\nabla f(x_{k})$, and Nozawa et al. [40] consider variants of (3) for problems with nonlinear inequality constraints. Grishchenko et al. [23] effectively consider the iteration (3) within a proximal point framework to handle an additional convex nonsmooth term in (1).
@@
Although not precisely based on sketching, many randomized algorithms have been proposed based on a related idea thoroughly expounded in [39] that was introduced as early as in [42][Chapter 3.4]. In such algorithms, one effectively draws one-dimensional sketches $S_{k}$ from a zero-mean, typically Gaussian, distribution with arbitrary covariance and then computes an unbiased estimator of $\nabla f(x_{k})$, $(\nabla S_{k}^{\top}\nabla f(x_{k}))S_{k}$. Randomized methods depending on a forward-mode AD oracle for computing $S_{k}^{\top}\nabla f(x_{k})$ in this unbiased estimator have also attracted recent attention in the AD community [5, 45], with a particular interest in application to fine-tuning of large language models [33, 56]. Nesterov and Spokoiny [39] also introduced a finite-difference-based (and hence biased) estimator for the derivative-free setting, which has inspired much research; see [6] and references therein.
@@
Subspace methods based on sketching have also inspired research in DFO. If $f$ is not readily amenable to AD, then one can approximate a given $S_{k}^{\top}\nabla f(x_{k})$ by finite differencing at the cost of $\mathcal{O}(p_{k})$ many function evaluations. This idea of approximating local projected gradient information lends itself naturally to model-based DFO methods, and this idea has been explored thoroughly in [12] and the previously mentioned [30]. Related analyses of subspace model-based DFO were performed in [13, 25]. Subspace model-based DFO was extended to stochastic objectives in [18].
@@
While the works cited so far concern randomized methods (i.e., $S_{k}$ is drawn from some distribution), we remind the reader that our broad definition in Framework 1 also encompasses deterministic methods. Block-coordinate descent methods qre widely found in the literature (see in particular [38, 51, 43]) and need not be randomized. A vast literature on conjugate gradient methods also exists; but we draw special attention to [55], which proposed a method that iteratively minimized a model of the objective restricted to a two-dimensional search space defined by the span of $\nabla f(x_{k})$ and the most recent displacement vector between incumbents. In our notation, this is tantamount to including $\nabla f(x_{k})$ as a column of $S_{k}$. Obviously, proposing to use such a sequence of $\{S_{k}\}$ in our setting, where we assume the full computation of $\nabla f(x_{k})$ is expensive, is nonsensical but would satisfy (2), which is the overarching goal of the UCB method we propose using in this manuscript. The identification of many well-known minimization methods effectively amounting to iterative minimization of local models along gradient-aligned subspaces is the subject of [53, 54]; this idea has permeated much development across the field. We also note that the Ph.D. thesis [57] considered, essentially, a derivative-free variant of the method of [55] that would periodically recompute a new approximation to $\nabla f(x_{k})$ for use in determining a subspace.
@@
We comment that the linear UCB mechanism we propose in this manuscript is meant to locally predict subspaces exhibiting the most variation, which are, in many applications, likely to change as the incumbent $x_{k}$ changes. We must mention that for problems that exhibit global subspaces of variation, there has been much research in identifying such so-called active subspaces [17]. In fact, some recent work has made explicit connections between sketching algorithms and identifying such global active subspaces [11, 10]. While our linear UCB mechanism is intuitively also likely to capture such low effective dimensionality, provided it exists for a given instance of $f$ in (1), our linear UCB mechanism is intended to identify, dynamically, local subspaces of variation. Nonetheless, in our numerical experiments we will test the hypothesis that the linear UCB mechanism should identify active subspaces.

### v4

@@ 1 Introduction
$\min_{\mathbf{x}\in\mathbb{R}^{d}}f(\mathbf{x}).$ (1)
Throughout the manuscript, we make the following assumption on $f$.
@@
  • $f:\mathbb{R}^{d}\to\mathbb{R}$ is continuously differentiable on a level set
    $L(\mathbf{x}_{0}):=\{\mathbf{z}\in\mathbb{R}^{d}:f(\mathbf{z})\leq f(\mathbf{x}_{0})\}$
  • $f$ is bounded below on $L(\mathbf{x}_{0})$; that is, there exists $\mathbf{x}_{*}\in\mathbb{R}^{d}$ such that $f(\mathbf{x}_{*})\leq f(\mathbf{x})$ for all $\mathbf{x}\in L(\mathbf{x}_{0})$.
  • $\nabla f$ is globally Lipschitz with constant $L$ on the level set $L(\mathbf{x}_{0})$; that is,
    $\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\|\leq L\|\mathbf{x}-\mathbf{y}\|$
    for all $\mathbf{x},\mathbf{y}\in L(\mathbf{x}_{0})$.
@@
In some areas of computational science, the computation of derivatives of objective function $f$ may be prohibitive even when 1 is satisfied. This is traditionally the scope of derivative-free optimization (DFO) methods; see, for example, [2, 17, 36]. Alternatively, for such functions $f$ that are provided as computer codes, algorithmic differentiation (AD) can provide numerically accurate derivative computations. A celebrated result (see, e.g., [25, Section 3.2]) proves that, in the absence of constraints on memory, reverse-mode AD can compute a numerical gradient of a function in a small factor times the wall-clock time of performing a zeroth-order function evaluation. Of course, these memory assumptions are unrealistic; and in practice, the computation of derivatives via reverse-mode AD is challenging, if not prohibitive. For a more thorough discussion of such challenges, see, for example, [30, Section 1]. In situations where reverse-mode AD is too difficult, however, forward-mode AD is sometimes tractable. Informally, forward-mode AD computes directional derivatives (in potentially arbitrary, that is, not coordinate-aligned, directions; see, for instance the jvp routine of JAX[8], which permits arbitrary tangent vectors.). The cost of a single directional derivative in this mode is approximately the cost of a single zeroth-order evaluation of $f$. This is, of course, an informal statement; see [30] for practical insight into more precise quantification of the true costs of forward-mode AD. Thus, one may view forward-mode AD as an oracle for computing directional derivatives (in arbitrary directions) at a cost similar to that of performing finite difference approximations of derivatives but with generally superior accuracy.
In computational settings that preclude analytic derivatives and efficient reverse-mode AD, model-based subspace optimization methods are of particular interest for the solution of (1), especially when the dimension $d$ is large, and have seen recent theoretical growth. For the sake of exposition, we broadly define model-based subspace methods as any iterative method of minimization that can be described by the informal statement provided in Framework 1.
1Input: Initial point $\mathbf{x}_{0}\in\mathbb{R}^{d}$, number of iterations $K$, function $f$ satisfying 1.
@@
3   Choose a subspace dimension $p_{k}\leq d$, and choose a matrix $\mathbf{S}_{k}\in\mathbb{R}^{d\times p_{k}}$.
4   Construct some model, $m_{k}:\mathbb{R}^{p_{k}}\to\mathbb{R}$, such that $m_{k}(\mathbf{z})$ approximates $f(\mathbf{x}_{k}+\mathbf{S}_{k}\mathbf{z})$ for all $\mathbf{z}\in\mathbb{R}^{p_{k}}$ in a neighborhood of $\mathbf{0}\in\mathbb{R}^{p_{k}}$.
5   Approximately minimize $m_{k}$ in a neighborhood of $\mathbf{0}\in\mathbb{R}^{p_{k}}$ to obtain $\mathbf{z}_{*}\in\mathbb{R}^{p_{k}}$.
6   Update the incumbent $\mathbf{x}_{k+1}\leftarrow\mathbf{x}_{k}+\mathbf{S}_{k}\mathbf{z}_{*}$, provided $f(\mathbf{x}_{k}+\mathbf{S}_{k}\mathbf{z}_{*})$ yields sufficient decrease.
@@
Framework 1 Informal definition of a model-based subspace optimization method
@@
Deterministic model-based subspace methods have been studied for decades; Krylov methods, and more specifically nonlinear conjugate gradient methods, fit within Framework 1 by implicitly choosing $\mathbf{S}_{k}$ to define a (typically) two-dimensional subspace based on past gradient evaluations. (L)BFGS methods implicitly generate $\mathbf{S}_{k}$ such that the span of the columns of $\mathbf{S}_{k}$ is equal to the span of a (limited) memory of gradients and past incumbent/gradient displacements. Block coordinate descent methods can be viewed as choosing $\mathbf{S}_{k}$ as a (potentially deterministic) subset of columns of a $d$-dimensional identity matrix. The former two types of subspace methods require the computation of (full-space) gradients, so these methods are not within the scope of our computational setting. Block coordinate descent methods do fit in our computational setting but are not the topic of this paper.
@@
Recent developments have seen the application of randomized matrix sketching [37, 58, 39, 43] as a means to choose $\mathbf{S}_{k}$ within algorithms in Framework 1, effectively choosing $\mathbf{S}_{k}$ as a sketch matrix, a random matrix governed by a distribution such that statistical guarantees can be made about how $\mathbf{S}_{k}$ acts on a vector (here, $\nabla f(\mathbf{x}_{k})$). In a computational setting that permits forward-mode AD, the local model $m_{k}$ of Framework 1 could be simply a first-order Taylor model in the affine subspace defined by $\mathbf{x}_{k}$ and $\mathbf{S}_{k}$. Computing $\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})$ amounts to computing a (relatively small) number of directional derivatives in the randomized directions given in the columns of $\mathbf{S}_{k}$. In the derivative-free setting, an arbitrary model $m_{k}$ is constructed in the (relatively small) affine subspace defined by $\mathbf{x}_{k}$ and $\mathbf{S}_{k}$; $m_{k}$ might exhibit the property that the zeroth- and first-order errors made by the model are comparable to those made by a first-order Taylor model. In either the forward-mode AD setting or the derivative-free setting, and as we will expound upon, the theoretical success of randomized subspace methods hinges upon some notion that the randomly selected subspace in each iteration exhibits, with sufficiently high probability, a sufficiently large normalized inner product with the gradient at the incumbent see, e.g., [9, Assumption 4][12, Assumptions 5,6][14, Assumption 5, Lemma 4][21, Assumptions 2,3]. All of these works assume, roughly speaking, that with high probability and for some $q\in(0,1)$, that
$\frac{\|(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\|}{\|\nabla f(\mathbf{x}_{k})\|}>q;$ (2)
Here, we have made the expositional choice to assume that $\mathbf{S}_{k}$ has full column rank so that a matrix with $p_{k}$ many columns always defines a $p_{k}$-dimensional subspace. Hence, $\mathbf{S}_{k}^{\top}\mathbf{S}_{k}$ is always invertible, and so the normalization by the matrix inverse in (2) is well-defined. One may notice that the normalization does not explicitly appear in the previously cited works, but is instead handled by assuming a (probabilistic) bound on $\|\mathbf{S}_{k}\|$.
@@
Providing a guarantee like (2) typically relies on some variant of the Johnson–Lindenstrauss lemma [31], a surprising result that demonstrates, roughly speaking, that in high dimensions, a low-dimensional subspace defined by $\mathbf{S}_{k}$ (with $p_{k}$ independent of $d$) drawn from an appropriate distribution achieves a property resembling (2) with high probability.
@@ 1.1 Contributions
In this manuscript we propose a novel deterministic method for choosing subspaces $\mathbf{S}_{k}$ within subspace optimization methods like those described by Framework 1. Instead of directly employing randomized sketches in an effort to (probabilistically) achieve a condition like (2), we demonstrate how achieving (2) can be performed by linear bandit methods in a nonstationary environment. Because 1 guarantees that the gradient of $f$ is Lipschitz continuous, we treat the incumbent update in a generic subspace optimization method as an exogenous process resulting in bounded drifts in a gradient.
@@
More specifically, we adapt a linear upper confidence bound (UCB) bandit method [3, 1] to choose each unit-length one-dimensional sketch $\mathbf{S}_{k}\in\mathbb{R}^{d}$ so that $\mathbf{S}_{k}$ is maximally aligned with $\nabla f(\mathbf{x}_{k})$; that is, we aim to maximize $|(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})|$ in every iteration. The linear UCB method maintains a confidence ellipsoid describing a belief on the dynamically changing value of $\nabla f(\mathbf{x}_{k})$ by fitting a linear regression model using a history of observed (approximations to) $\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})$. The maximizer over this confidence ellipsoid of an inner product with a unit-norm vector is chosen on each iteration as the one-dimensional sketch $\mathbf{S}_{k}$. Such a UCB method ensures a balance between exploration of subspaces—the focus of randomized sketching approaches to subspace optimization methods in the literature—and exploitation of subspaces that have recently been observed to exhibit large gradients. The linear UCB method is introduced and analyzed in Section 3.
@@
In Section 3 we also demonstrate sublinear dynamic regret [6, 7] of this linear UCB method in terms of its ability to approximate the dynamically changing gradient $\nabla f(\mathbf{x}_{k})$. Because of our choice to view the optimization dynamics as an exogenous process affecting the gradient $\nabla f(\mathbf{x}_{k})$, we stress that this is not a regret bound for the optimization algorithm. However, our regret analysis will elucidate how the selection of a particular subspace optimization algorithm will affect the regret incurred by the linear UCB method. As an illustration, we will consider perhaps the simplest possible optimization algorithm that could fit into Framework 1—a subspace variant of gradient descent with a backtracking line search—to make these connections clear. In particular, given a sequence $\{\mathbf{S}_{k}\}$ of full column rank matrices, define the associated sequence of embedding matrices $\{\mathbf{P}_{k}:=\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}\}\subset\mathbb{R}^{d\times d}$. Then, we consider the following algorithm:
1
2(Inputs) Initial point $\mathbf{x}_{1}$, horizon $K$.
3(Initializations) Backtracking parameter $\beta\in(\frac{1}{2},1)$.
4for $k=1,2,\dots,K$ do
5 (Obtain (random) sketch)
6  Receive $\mathbf{S}_{k}\in\mathbb{R}^{d\times p_{k}}$ and call oracle to compute $\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})$.
7 (Determine next incumbent by linesearch)
8 $\alpha\leftarrow 1.$
9 $\mathbf{x}_{+}\leftarrow\mathbf{x}_{k}-\alpha\mathbf{P}_{k}\nabla f(\mathbf{x}_{k})$.
10 while $f(\mathbf{x}_{+})\geq f(\mathbf{x}_{k})-\alpha\|\mathbf{P}_{k}\nabla f(\mathbf{x}_{k})\|^{2}$ do
11
12    $\alpha\leftarrow\beta\alpha$
13    $\mathbf{x}_{+}\leftarrow\mathbf{x}_{k}-\alpha\mathbf{P}_{k}\nabla f(\mathbf{x}_{k})$.
14
15 $\mathbf{x}_{k+1}\leftarrow\mathbf{x}_{+}.$
Algorithm 2 Subspace Gradient Descent with Backtracking Line Search.
Algorithm 2 is among the simplest algorithms that could fit into Framework 1 because the subspace model $m_{k}$ employed in each iteration is a first-order Taylor model restricted to the subspace defined by $\mathbf{S}_{k}$, and a quantifiable notion of sufficient decrease from $m_{k}$ is classically guaranteed by the backtracking line search for $f$ satisfying 1.
@@ 12    $\alpha\leftarrow\beta\alpha$
We are particularly concerned in this manuscript with practicalities; and so, motivated by various pieces of the analysis in Section 3, we propose practical extensions in Section 4. We conclude this paper by proposing and testing our linear UCB mechanism in the practical variant of Algorithm 2 in a setting where we simulate the computation of forward-mode AD gradients. In a derivative-free setting, we propose a simple extension of the DFO method POUNDers [56] that employs our linear UCB mechanism to select subspaces.
@@ 1.2 Related Work
The theoretical and practical development of methods that have the general form of Framework 1 has enjoyed much recent attention. Matrix sketching [37, 58, 39, 43], which entails the random generation of $\mathbf{S}_{k}$ in Framework 1 from some distribution on matrices, has emerged as a theoretical tool for analyzing such methods, often supported by the powerful Johnson–Lindenstrauss lemma [31]. In [34] the authors propose a method in Framework 1 that looks much like the subspace gradient descent method in Algorithm 2, but with a fixed step size. The authors prove convergence of the method under standard assumptions about convexity of $f$; in a follow-up paper [35], the same authors considered the same iteration but with $\nabla f(\mathbf{x}_{k})$ approximated by finite differences. The works [24, 60, 22] and [28, 66] propose and analyze randomized Newton methods and stochastic subspace cubic Newton methods, respectively, that extend a subspace gradient descent iteration by also using second-derivative information projected onto subspaces defined by $S_{k}$. Pilanci and Wainwright [47] also similarly consider randomized Newton methods in the case where the objective Hessian admits a matrix square root everywhere on its domain. Cartis et al. [9] consider imposing various globalization strategies, such as trust-region and quadratic regularization, into algorithms that fall into Framework 1; the authors also specialize their methods to nonlinear least-squares objectives. Miyashi et al. [41] consider a quasi-Newton approximation of projected Hessian information to be used with projected gradients $\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})$, and Nozawa et al. [46] consider variants of the subspace gradient descent iteration for problems with nonlinear inequality constraints. Grishchenko et al. [26] effectively consider the subspace gradient descent iteration within a proximal point framework to handle an additional convex nonsmooth term in (1).
@@
Although not precisely based on sketching, many randomized algorithms have been proposed based on a related idea thoroughly expounded in [45] that was introduced as early as in [48, Chapter 3.4]. In such algorithms, one effectively draws one-dimensional sketches $\mathbf{S}_{k}\in\mathbb{R}^{d}$ from a zero-mean, typically Gaussian, distribution with arbitrary covariance and then computes an unbiased estimator $\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k}))\mathbf{S}_{k}$ of $\nabla f(\mathbf{x}_{k})$. Randomized methods depending on a forward-mode AD oracle for computing $\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})$ in this unbiased estimator have also attracted recent attention in the AD community [4, 52], with a particular interest in application to fine-tuning of large language models [38, 64].
@@
Model-based subspace methods based on sketching have also inspired research in DFO. If $f$ is not readily amenable to AD, then one can approximate a given $\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})$ by finite differencing at the cost of $\mathcal{O}(p_{k})$ many function evaluations. This idea of approximating local projected gradient information lends itself naturally to model-based DFO methods, and this idea has been explored thoroughly in [12, 13] and the previously mentioned [35]. Related analyses of subspace model-based DFO were performed in [14, 29]. Subspace model-based DFO was extended to stochastic objectives in [21]and to more general noisy objectives in [20]. Although not model-based methods, randomized subspace techniques have also been explored in the context of direct-search methods, another well-studied class of methods in DFO [50, 19]. We additionally mention that Nesterov and Spokoiny [45] also introduced a finite-difference-based (and hence biased) estimator for the derivative-free setting, which has inspired much research; see [5] and references therein.
@@
While the works cited so far concern randomized methods (i.e., $\mathbf{S}_{k}$ is drawn from some distribution), we remind the reader that our broad definition in Framework 1 also encompasses deterministic methods. Block-coordinate descent methods are widely found in the literature (see in particular [44, 59, 49]) and need not be randomized. A vast literature on conjugate gradient methods also exists; but we draw special attention to [63], which proposed a method that iteratively minimized a model of the objective restricted to a two-dimensional search space defined by the span of $\nabla f(\mathbf{x}_{k})$ and the most recent displacement vector between incumbents. In our notation, this is tantamount to including $\nabla f(\mathbf{x}_{k})$ as a column of $\mathbf{S}_{k}$. Obviously, proposing to use such a sequence of $\{\mathbf{S}_{k}\}$ in our setting, where we assume the full computation of $\nabla f(\mathbf{x}_{k})$ is expensive, is nonsensical but would satisfy (2), which is the overarching goal of the UCB method we propose using in this manuscript. The identification of many well-known minimization methods effectively amounting to iterative minimization of local models along gradient-aligned subspaces is the subject of [61, 62]; this idea has permeated much development across the field. We also note that the Ph.D. thesis [65] considered, essentially, a derivative-free variant of the method of [63] that would periodically recompute a new approximation to $\nabla f(\mathbf{x}_{k})$ for use in determining a subspace.
@@
We comment that the linear UCB mechanism we propose in this manuscript is meant to locally predict subspaces exhibiting the most variation, which are, in many applications, likely to change as the incumbent $\mathbf{x}_{k}$ changes; as will be seen in Algorithm 4, an employed memory parameter $M$ will guarantee that the UCB method only retains the $M$ most recent sketched gradients in memory, justifying this locality. We must mention that for problems that exhibit global subspaces of variation, there has been much research in identifying such so-called active subspaces [18]. In fact, some recent work has made explicit connections between sketching algorithms and identifying such global active subspaces [11, 10]. While our linear UCB mechanism is intuitively also likely to capture such low effective dimensionality, provided it exists for a given instance of $f$ in (1), our linear UCB mechanism is intended to identify, dynamically, local subspaces of variation. Nonetheless, in our numerical experiments we will test the hypothesis that the linear UCB mechanism should identify active subspaces.